The formula A = 6V^2/3 gives the surface area of a cube in terms of its volume V. What is the volume of a cube with surface area 12 square feet?
Answer by checkley75(3666)

A=6*V^2/3
12=6V^2/3
12/6=V^2/3
2=V^2/3 MULTIPLYING BOTH SIDES BY (^3/2) WE GET
V^(2/3)(3/2)=2^3/2
V^1=2^3/2
V=SQRT2^3
V=2SQRT2